The following information has been taken from the accounting records of Klear-Seal Company for last year.

Selling expenses                                               140,000

Raw materials inventory, January 1                      90,000

Raw materials inventory, December 31                60,000

Utilities, factory                                                    36,000

Direct labour cost                                              150,000

Depreciation, factory                                         162,000

Purchases of raw materials                               750,000

Sales                                                               2,500,000

Insurance, factory                                                 40,000

Supplies, factory                                                   15,000

Administrative expenses                                    270,000

Indirect labour                                                   300,000

Maintenance, factory                                            87,000

Work in process inventory, January 1                180,000

Work in process inventory, December 31          100,000

Finished goods inventory, January 1                 260,000

Finished goods inventory, December 31                       210,000

Management wants to organize these data into a better format so that financial statements can be prepared for the year.

Required:

  1. Prepare a schedule of cost of goods manufactured.

  1. Compute the cost of goods sold.
  1. Using data as needed from (1) and (2) above, prepare an income statement.
  1. Assuming production of finished and semi-finished goods amounted to 412,500 units for the past year. Calculate the cost components of the ending finished goods inventory of 55,176 units. (Hint: The categories of costs in the ending inventory are the same as the total manufacturing costs; only the amounts are different.)
Answer #1
Particulars Amount Amount
Raw Materials Inventory Jan 01 90,000
Add : Purchases of Raw Material 750,000
Raw Materials available for use 840,000
Less: Raw Materials Inventory Dec 31 60,000
Raw Materials Consumed 780,000
Direct Labour 150,000
Factory Overhead
Utilities 36,000
Depreciation 162,000
Insurance 40,000
Supplies 15,000
Indirect Labour 300,000
Maintenance Overhead 87,000
Total Factory Overheads 640,000
Total Manufacturing Costs 1,570,000
Add: Work in Process Inventory January 1 180,000
Total Work in Process 1,750,000
Less: Work in Process Inventory Dec 31 100,000
Cost of Goods Manufactured 1,650,000

Total Manufacturing Costs = Direct Materials Used + Direct Labour Cost + Total Manufacturing Overhead

Cost of Goods Sold Statement

Particulars Amount Amount
Cost of Goods Manufactured 1,650,000
Add Finished Goods Inventory Jan 01 260,000
Cost of Goods available for Sale 1,910,000
Less Finished Goods Dec31 210,000
Cos of Goods Sold 1,700,000

Income Statement

Particulars Amount Amount
Sales 2,500,000
Less: Cost of Goods Sold 1,700,000
Gross Profit 800,000
Selling and Administrative Expenses
Selling Expenses 140,000
Administrative Expense 270,000
Total Selling and Administrative Expenses 410,000
Operating Income 390,000

Break Down of Finished Goods Inventory of 55,176 Units

Raw Materials Costs = Total Cost of Raw Materials Used / 412,500 Units

Raw Materials Cost per Unit = 780,000 / 412,500 = 1.89 Per Unit

Direct Labour Cost per Unit = 150,000 / 412,500 = $ 0.36 per Unit

Manufacturing Overhead per Unit = 640,000 / 412,500 = $ 1.55 per Unit

Ending Inventory

A B C = A*B
Particulars Units Per Unit Total Cost
Direct Materials 55,176 1.89 104,333
Direct Labour Cost 55,176 0.36 20,064
Manufacturing Overhead 55,176 1.55 85,606
Total Manufacturing Costs 210,003

In Calculation of Decomposition of Ending Inventory onto Various Components Calculation regarding most of the items were in fraction so rounding off has been done.